# Note on a rigorous derivation of self-consistent double-hybrid functional theory via generalized Kohn-Sham theory and cumulant approximation

### 12 Mar 2026 v1

@@ Abstract
In this short note, we present a rigorous theoretical derivation of the one-body double-hybrid density functional (OBDHF) theory, a novel self-consistent double-hybrid density functional framework that unifies the generalized Kohn-Sham (GKS) formalism with one-body Møller-Plesset second-order perturbation (OBMP2) theory. Conventional double-hybrid density functionals suffer from a fundamental theoretical inconsistency arising from the non-self-consistent treatment of the perturbative MP2 correlation, in which the orbitals entering the correlation energy expression are not variationally optimized with respect to the full double-hybrid energy functional. To address this deficiency, we construct a model energy functional as a linear combination of semilocal density functional approximation XC, a fraction $\alpha_{x}$ of exact Hartree-Fock (HF) exchange, and a fraction $\alpha_{c}$ of OBMP2 correlation. By virtue of the one-body operator structure of OBMP2, the perturbative correlation contribution is embedded directly and self-consistently into the GKS effective Hamiltonian, without recourse to the optimized effective potential (OEP) construction or perturbative orbital relaxation corrections. Through functional differentiation of the total OBDHF energy with respect to the orbitals, we derive the OBDHF effective Hamiltonian and the associated self-consistent field equations in a rigorous and transparent manner. This formulation provides a theoretically well-founded and practically tractable pathway to fully self-consistent double-hybrid density functional theory calculations within the GKS framework, resolving the self-consistency problem inherent in conventional double-hybrid functionals.
@@ I Introduction
Despite their remarkable accuracy, conventional DH functionals suffer from a fundamental theoretical inconsistency: the MP2 correlation energy is evaluated as a non-self-consistent, post-SCF correction applied to orbitals optimized under the hybrid XC functional without the MP2 contribution. As a consequence, the orbitals entering the perturbative correlation expression are not variationally optimized with respect to the full DH energy functional. This non-self-consistency implies that the one-particle density matrix, and hence physical observables sensitive to it, such as the electron density, dipole moment, and response properties, are not fully consistent with the DH energy expression Sharkas et al. (2011); Toulouse et al. (2011); Fromager (2011). Achieving a genuinely self-consistent treatment of DH functionals therefore represents an important open challenge in the development of orbital-dependent density functional approximations.
@@
A particularly natural framework for embedding MP2-level correlation self-consistently into a GKS orbital equation is offered by the one-body Møller-Plesset second-order perturbation (OBMP2) theory, recently developed by Tran and coworkers Tran and Yanai (2013); Tran (2021, 2022); Le and Tran (2023); Tran et al. (2024); Tran and Tran (2025). OBMP2 reformulates the MP2 dynamic correlation effects as an effective one-body correlated Fock operator, obtained through a unitary canonical transformation of the molecular Hamiltonian Yanai and Chan (2006, 2007); Chan and Yanai (2007); Neuscamman et al. (2009, 2010b, 2010a) truncated at the second order. Because OBMP2 yields a one-body operator rather than a two-body perturbative energy, it is uniquely suited for direct incorporation into the GKS effective Hamiltonian without recourse to response equations, orbital gradient corrections, or the OEP construction.
In this note, we present the development of self-consistent one-body double-hybrid (OBDHF) density functional theory that combines the GKS framework with OBMP2 theory. The central contribution of this note is a rigorous derivation of the OBDHF orbital equation, in which both a fraction $\alpha_{x}$ of exact HF exchange and a fraction $\alpha_{c}$ of OBMP2 correlation are embedded self-consistently within the GKS model functional and the corresponding orbital optimization. The model functional is constructed as a linear combination of semilocal DFA XC, exact exchange, and OBMP2 correlation contributions. By taking the functional derivative of the total energy with respect to the orbitals, we derive the OBDHF effective Hamiltonian and the associated self-consistent field (SCF) equations. This formulation provides, for the first time, a theoretically rigorous and practically tractable pathway to fully self-consistent double-hybrid DFT calculations within the GKS framework, without invoking the OEP or perturbative orbital relaxation corrections.
@@ II.1 Generalized Kohn-Sham theory
$E_{tot}=\min_{n}\left(\int V_{ext}(\mathbf{r})n(\mathbf{r})d\mathbf{r}+F_{S}[n]+R_{S}[n]\right)$ (6)
@@
$\left(\hat{O}_{S}[\{\phi_{j}\}]+V_{ext}(\mathbf{r})+V_{R}(\mathbf{r})\right)\phi_{i}(\mathbf{r})=\varepsilon_{i}\phi_{i}(\mathbf{r})$ (7)
This equation contains three distinct terms:
  • $V_{ext}(\mathbf{r})$: The external ionic potential.
@@
    $V_{R}(\mathbf{r})\equiv\frac{\delta R_{S}[n]}{\delta n(\mathbf{r})}$ (8)
@@ II.3 Self-consistent one-body double-hybrid functional (OBDHF) theory
In this section, we present a novel self-consistent OBDHF theory that combines the GKS framework with the OBMP2 theory. The central idea is to construct a model functional $S[\{\phi_{j}\}]$ that incorporates both a fraction of exact (HF) exchange and a fraction of OBMP2 correlation energy, embedding these contributions self-consistently within the GKS orbital equation.
II.3.1 Definition of the model functional
We define the model energy functional as a linear combination of HF-type kinetic and exchange contributions, a semi-local XC functional, and an OBMP2 correlation term:
@@ II.3.1 Definition of the model functional
  • $E_{\mathrm{H}}[n]$ is the classical Hartree (Coulomb) energy.
  • $E_{x}^{\mathrm{DFA}}[n]$ and $E_{c}^{\mathrm{DFA}}[n]$ are the exchange and correlation energies from a chosen density functional approximation (DFA).
@@
  • $\alpha_{x}\in[0,1]$ and $\alpha_{c}\in[0,1]$ are the mixing parameters for exact exchange and OBMP2 correlation, respectively.
II.3.2 OBMP2 correlation energy
The OBMP2 correlation energy is defined as the expectation value of $\hat{v}^{\mathrm{OBMP2}}$ (Eq. 18) with respect to the reference Slater determinant $|\Phi_{0}\rangle$:
@@ II.3.2 OBMP2 correlation energy
To evaluate Eq. 23, we use the standard result for the expectation value of a one-body operator $\hat{a}^{p}_{q}=\hat{a}^{\dagger}_{p}\hat{a}_{q}$ with respect to a single Slater determinant:
$\langle\Phi_{0}|\hat{a}^{p}_{q}|\Phi_{0}\rangle=\delta_{pq}\,n_{p},$ (24)
This means that $\langle\hat{\Omega}(\hat{a}^{p}_{q})\rangle$ is nonzero only when $p=q$ and the orbital is occupied. For all off-diagonal terms ($p\neq q$) or for virtual-virtual and occupied-virtual pairs with $p\neq q$, the expectation value vanishes identically. After applying Eq. 25 to all $\hat{\Omega}$ terms and retaining only the occupied-diagonal contributions, the surviving terms are:
Applying the Kronecker deltas and using $n_{j}=1$ for occupied orbitals, each surviving $\hat{\Omega}$ term collapses. For example, $g^{ip}_{ab}\cdot 2\delta_{pj}n_{j}=2g^{ij}_{ab}$, and noting that $\bar{T}^{ab}_{ij}\cdot g^{ij}_{ab}=T^{ab}_{ij}g^{ij}_{ab}-T^{ab}_{ji}g^{ij}_{ab}$. After collecting all terms and using the antisymmetry of $\bar{T}^{ab}_{ij}$ and $g^{ab}_{ij}$, the occ-occ diagonal $\hat{\Omega}$ contributions exactly cancel with part of the scalar terms. The final expression for the OBMP2 correlation energy is:
$E_{c}^{\mathrm{OBMP2}}[\{\phi_{j}\}]=\langle\Phi_{0}|\hat{v}^{\mathrm{OBMP2}}|\Phi_{0}\rangle=\bar{T}^{ab}_{ij}\,g^{ij}_{ab}\;+\;2\,f^{a}_{c}\,T^{ab}_{ij}\,T^{cb}_{ij}\;-\;2\,f^{k}_{i}\,T^{ab}_{ij}\,T^{ab}_{kj},$ (27)
where Einstein’s summation convention is used over all repeated indices, the antisymmetrized amplitude $\bar{T}^{ab}_{ij}=T^{ab}_{ij}-T^{ab}_{ji}$ is defined below Eq. 18, the MP2 amplitude $T^{ab}_{ij}$ is given in Eq. 12, $g^{ij}_{ab}$ are two-electron integrals, and $f^{p}_{q}$ is the Fock matrix of Eq. 19.
Several important observations follow from Eq. 27:
  • Reduction to standard MP2 in the canonical HF limit. When the orbitals are canonical HF orbitals, the Fock matrix is diagonal: $f^{p}_{q}=\varepsilon_{p}\,\delta_{pq}$. The virtual-virtual block gives $f^{a}_{c}=\varepsilon_{a}\,\delta_{ac}$, and the occupied-occupied block gives $f^{k}_{i}=\varepsilon_{i}\,\delta_{ki}$. Substituting into Eq. 27:
    $2f^{a}_{c}T^{ab}_{ij}T^{cb}_{ij}$ $=2\varepsilon_{a}T^{ab}_{ij}T^{ab}_{ij},$
    $2f^{k}_{i}T^{ab}_{ij}T^{ab}_{kj}$ $=2\varepsilon_{i}T^{ab}_{ij}T^{ab}_{ij},$
    so the last two terms combine as $2(\varepsilon_{a}-\varepsilon_{i})T^{ab}_{ij}T^{ab}_{ij}$. Together with the first term, substituting $T^{ab}_{ij}=g^{ab}_{ij}/(\varepsilon_{i}+\varepsilon_{j}-\varepsilon_{a}-\varepsilon_{b})$, one recovers the standard MP2 correlation energy:
    $E_{c}^{\mathrm{MP2}}=\bar{T}^{ab}_{ij}\,g^{ij}_{ab}=\sum_{ij}^{\mathrm{occ}}\sum_{ab}^{\mathrm{vir}}\frac{g^{ab}_{ij}\left(g^{ij}_{ab}-g^{ij}_{ba}\right)}{\varepsilon_{i}+\varepsilon_{j}-\varepsilon_{a}-\varepsilon_{b}}.$ (28)
  • Orbital-relaxation corrections. In the self-consistent OBDHF procedure, the orbitals are no longer canonical HF orbitals, and $f^{p}_{q}$ acquires off-diagonal elements. The terms $2f^{a}_{c}T^{ab}_{ij}T^{cb}_{ij}$ and $2f^{k}_{i}T^{ab}_{ij}T^{ab}_{kj}$ then represent orbital-relaxation corrections beyond standard MP2, arising from the back-polarization of the orbitals by the correlated potential $\hat{v}^{\mathrm{OBMP2}}$.
II.3.3 Remainder energy functional
@@
which absorbs all many-body contributions not captured by $S[\{\phi_{j}\}]$. The corresponding multiplicative remainder potential is:
@@
II.3.4 The OBDHF orbital equation
Substituting the model functional $S[\{\phi_{j}\}]$ into the GKS equation (7), we obtain the self-consistent OBDHF orbital equation:
@@ II.3.4 The OBDHF orbital equation
where:
  • $V_{\mathrm{H}}(\mathbf{r})=\int\frac{n(\mathbf{r}^{\prime})}{|\mathbf{r}-\mathbf{r}^{\prime}|}d\mathbf{r}^{\prime}$ is the Hartree potential (Eq. 3).
  • $\hat{v}_{x}^{\mathrm{HF}}[\{\phi_{j}\}]$ is the non-multiplicative HF exchange operator:
@@
  • $\hat{v}^{\mathrm{OBMP2}}[\{\phi_{j}\}]$ is the non-multiplicative OBMP2 correlated potential operator derived from $\hat{H}^{\mathrm{OBMP2}}$ (Eq. 1718), whose matrix elements in the spin-orbital basis are given by:
    $v_{q}^{p}=\bar{f}_{q}^{p}-f_{q}^{p},$ (33)
    with the correlated Fock matrix $\bar{f}_{q}^{p}$ defined in Eq. 21.
  • $V_{R}(\mathbf{r})$ is the multiplicative remainder potential defined in Eq. 30.
The non-multiplicative operator $\hat{O}_{S}[\{\phi_{j}\}]$ in the general GKS Eq. 7 is here explicitly identified as:
@@ II.3.5 Total energy expression
II.3.6 Self-consistent field procedure
@@
Step 2. Amplitude update: Compute the MP2 amplitudes $T_{ij}^{ab}$ (Eq. 12) using the current orbital energies $\{\varepsilon_{i}\}$.
Step 3. Correlated potential update: Construct the OBMP2 correlated potential $\hat{v}^{\mathrm{OBMP2}}[\{\phi_{j}\}]$ from Eqs. 1721 and 33.
Step 4. Fock matrix assembly: Assemble the effective OBDHF Fock matrix:
@@
Step 5. Orbital update: Diagonalize $F^{\mathrm{OBDHF}}$ to obtain new orbitals $\{\phi_{i}^{\mathrm{new}}\}$ and eigenvalues $\{\varepsilon_{i}^{\mathrm{new}}\}$.
Step 6. Convergence check: Repeat steps 2–5 until the change in the total energy and the density matrix fall below the desired threshold:
@@
Upon convergence, the self-consistent orbitals incorporate both the non-local HF exchange and the dynamic correlation encoded in $\hat{v}^{\mathrm{OBMP2}}$ on an equal footing, going beyond the standard post-SCF double-hybrid approach. The resulting orbital set minimizes the total energy functional (35) within the GKS framework, ensuring a fully variational and self-consistent treatment of the double-hybrid functional.
@@ III Discussion and Outlook
In this note, we have presented a rigorous theoretical derivation of the one-body double-hybrid density functional (OBDHF) theory, a self-consistent double-hybrid density functional framework grounded in the generalized Kohn-Sham (GKS) formalism and the one-body Møller-Plesset second-order perturbation (OBMP2) theory. The central theoretical challenge addressed here is the long-standing self-consistency problem inherent in conventional double-hybrid density functionals, wherein the perturbative MP2 correlation energy is evaluated as a non-self-consistent, post-SCF correction on top of orbitals that are not variationally optimized with respect to the full double-hybrid energy functional.
The key insight exploited in the present development is the one-body operator structure of OBMP2, which encapsulates second-order dynamic correlation effects through an effective correlated Fock operator derived via a unitary canonical transformation of the molecular Hamiltonian. This one-body character renders OBMP2 uniquely amenable to direct and self-consistent embedding within the GKS effective Hamiltonian, in sharp contrast to the conventional two-body MP2 energy expression, which cannot be straightforwardly incorporated into a GKS orbital equation without invoking either the optimized effective potential construction or perturbative orbital relaxation corrections.
@@
The theoretical framework presented here thus provides, for the first time, a well-founded and practically tractable route to fully self-consistent double-hybrid DFT calculations within the GKS framework. The OBDHF formulation is expected to yield improved consistency between the energy and the one-particle density matrix relative to conventional double-hybrid functionals, with potential benefits for the computation of electron densities, dipole moments, and response properties. A comprehensive numerical implementation of the OBDHF equations, together with an extensive assessment of its performance across thermochemistry, kinetics, non-covalent interactions, and molecular properties benchmarks, is currently underway and will be reported in a forthcoming paper.

### 14 Mar 2026 v2

@@ Abstract
In this short note, we present a rigorous derivation of the one-body double-hybrid density functional (OBDHF) theory, a self-consistent double-hybrid density functional framework that unifies the generalized Kohn-Sham (GKS) formalism with one-body Møller-Plesset second-order perturbation (OBMP2) theory. Conventional double-hybrid density functionals suffer from a fundamental theoretical inconsistency arising from the non-self-consistent treatment of the perturbative MP2 correlation, in which the orbitals entering the correlation energy expression are not variationally optimized with respect to the full double-hybrid energy functional. To address this deficiency, we construct a model energy functional as a linear combination of semilocal density functional approximation XC, a fraction $\alpha_{x}$ of exact Hartree-Fock (HF) exchange, and a fraction $\alpha_{c}$ of OBMP2 correlation. By virtue of the one-body operator structure of OBMP2, the perturbative correlation contribution is embedded directly and self-consistently into the GKS effective Hamiltonian, without recourse to the optimized effective potential (OEP) construction or perturbative orbital relaxation corrections. Through functional differentiation of the total OBDHF energy with respect to the orbitals, we derive the OBDHF effective Hamiltonian and the associated self-consistent field equations in a rigorous and transparent manner. This formulation provides a theoretically well-founded and practically tractable pathway to fully self-consistent double-hybrid density functional theory calculations within the GKS framework, resolving the self-consistency problem inherent in conventional double-hybrid functionals.
@@ I Introduction
Despite their remarkable accuracy, conventional DH functionals suffer from a fundamental theoretical inconsistency: the MP2 correlation energy is evaluated as a non-self-consistent, post-SCF correction applied to orbitals optimized under the hybrid XC functional without the MP2 contribution. As a consequence, the orbitals entering the perturbative correlation expression are not variationally optimized with respect to the full DH energy functional. This non-self-consistency implies that the one-particle density matrix, and hence physical observables sensitive to it, such as the electron density, dipole moment, and response properties, are not fully consistent with the DH energy expression Sharkas et al. (2011); Toulouse et al. (2011); Fromager (2011). Achieving a self-consistent treatment of DH functionals therefore represents an important open challenge in the development of orbital-dependent density functional approximations.
@@
A particularly natural framework for wrapping MP2-level correlation self-consistently into a GKS orbital equation is offered by the one-body Møller-Plesset second-order perturbation (OBMP2) theory, recently developed by the author Tran and Yanai (2013); Tran (2021, 2022); Le and Tran (2023); Tran et al. (2024); Tran and Tran (2025). OBMP2 formulates the MP2 dynamic correlation effects as an effective one-body correlated Fock operator, obtained through a unitary canonical transformation of the molecular Hamiltonian Yanai and Chan (2006, 2007); Chan and Yanai (2007); Neuscamman et al. (2009, 2010b, 2010a) followed by the cummulant approximation to reduce many-body operators to one-body operators. Because OBMP2 yields a one-body correlated potential operator rather than a two-body perturbative energy, it is uniquely suited for direct incorporation into the GKS effective Hamiltonian without recourse to response equations, orbital gradient corrections, or the OEP construction.
In this note, we present the development of self-consistent one-body double-hybrid (OBDHF) density functional theory that combines the GKS framework with OBMP2 theory. The central contribution of this note is a rigorous derivation of the OBDHF orbital equation, in which both a fraction $\alpha_{x}$ of exact HF exchange and a fraction $\alpha_{c}$ of OBMP2 correlation are embedded self-consistently within the GKS model functional and the corresponding orbital optimization. The model functional is constructed as a linear combination of semilocal DFA XC, exact exchange, and OBMP2 correlation. By taking the functional derivative of the total energy with respect to the orbitals, we derive the OBDHF effective Hamiltonian and the associated self-consistent field (SCF) equations. This formulation provides a theoretically rigorous and practically tractable pathway to fully self-consistent double-hybrid DFT calculations within the GKS framework, without invoking the OEP or perturbative orbital relaxation corrections.
@@ II.1 Generalized Kohn-Sham theory
$E_{tot}=\min_{n}\left(\int V_{ext}(\mathbf{r})n(\mathbf{r})d\mathbf{r}+F_{S}[n]+R_{S}[n]\right).$ (6)
@@
$\left(\hat{O}_{S}[\{\phi_{j}\}]+V_{ext}(\mathbf{r})+V_{R}(\mathbf{r})\right)\phi_{i}(\mathbf{r})=\varepsilon_{i}\phi_{i}(\mathbf{r}),$ (7)
where:
@@
    $V_{R}(\mathbf{r})\equiv\frac{\delta R_{S}[n]}{\delta n(\mathbf{r})}.$ (8)
@@ II.3 Self-consistent one-body double-hybrid functional (OBDHF) theory
In this section, we present the self-consistent OBDHF theory combining the GKS framework with the OBMP2 theory. The idea is to construct a model functional $S[\{\phi_{j}\}]$ incorporating both a fraction of exact (HF) exchange and a fraction of OBMP2 correlation energy, including these contributions self-consistently within the GKS orbital equation. Let us first define the model energy functional as a linear combination of HF-type kinetic and exchange contributions, a semi-local XC functional, and an OBMP2 correlation term:
@@
  • $E_{\mathrm{H}}[n]$ is the classical Hartree energy.
  • $E_{x}^{\mathrm{DFA}}[n]$ and $E_{c}^{\mathrm{DFA}}[n]$ are the exchange and correlation energies of a density functional approximation (DFA).
@@
  • $\alpha_{x}\in[0,1]$ and $\alpha_{c}\in[0,1]$ are the mixing parameters for the exact exchange and the OBMP2 correlation, respectively.
The OBMP2 correlation energy is defined as the expectation value of $\hat{v}^{\mathrm{OBMP2}}$ (Eq. 18) with respect to the reference Slater determinant $|\Phi_{0}\rangle$ constituted by the orbital set $\{\phi_{j}\}$:
@@
consisting of all many-body effects not captured by $S[\{\phi_{j}\}]$. The corresponding multiplicative remainder potential is:
@@
Substituting the model functional $S[\{\phi_{j}\}]$ into the GKS equation (Eq. 7), we obtain the self-consistent OBDHF orbital equation:
@@
where $\hat{v}_{x}^{\mathrm{HF}}[\{\phi_{j}\}]$ is the HF exchange operator:
@@
The non-multiplicative operator $\hat{O}_{S}[\{\phi_{j}\}]$ in Eq. 7 is now explicitly defined as:
@@
II.4 OBDHF self-consistency
@@
Step 2. Updating amplitude: Calculate the MP2 amplitudes $T_{ij}^{ab}$ (Eq. 12) using the current orbitals and orbital energies $\{\varepsilon_{i}\}$.
Step 3. Updating correlated potential: Construct the OBMP2 correlated potential $\hat{v}^{\mathrm{OBMP2}}[\{\phi_{j}\}]$.
Step 4. Calculating Fock matrix: Construct the effective OBDHF Fock matrix:
@@
Step 5. Updating orbital: Diagonalize $F^{\mathrm{OBDHF}}$ to obtain new orbitals $\{\phi_{i}^{\mathrm{new}}\}$ and eigenvalues $\{\varepsilon_{i}^{\mathrm{new}}\}$.
Step 6. Check convergence: Repeat steps 2–5 until the change in the total energy (Eq. 29) and the density matrix fall below the desired threshold:
@@
Upon convergence, OBDHF orbitals incorporate both the non-local HF exchange and the dynamic correlation encoded in $\hat{v}^{\mathrm{OBMP2}}$ on an equal footing, going beyond the standard non-self-consistent DH approach. The resulting set of orbitals minimizes the total energy functional (29) within the GKS framework, ensuring a fully variational and self-consistent treatment of the DH functional.
@@ III Discussion and Outlook
In this note, we have presented a rigorous derivation of the OBDHF theory, a self-consistent double-hybrid density functional framework grounded in the generalized Kohn-Sham (GKS) formalism and the OBMP2 theory. The central theoretical challenge addressed here is the long-standing self-consistency problem inherent in conventional DH density functionals, wherein the perturbative MP2 correlation energy is evaluated as a non-self-consistent, post-SCF correction on top of KS orbitals that are not variationally optimized with respect to the full DH energy functional.
The key insight exploited in the present development is the one-body operator structure of OBMP2, which encapsulates second-order dynamic correlation effects through an effective correlated Fock operator derived via a unitary canonical transformation of the molecular Hamiltonian followed by the cummulant approximation to reduce high-order operators to one-body operators. This one-body character renders OBMP2 uniquely amenable to direct and self-consistent embedding within the GKS effective Hamiltonian, in contrast to the conventional two-body MP2 energy expression, which cannot be straightforwardly incorporated into a GKS orbital equation without invoking either the OEP construction or perturbative orbital relaxation corrections.
@@
The theoretical framework presented here thus provides a well-founded and practically tractable route to fully self-consistent double-hybrid DFT calculations within the GKS framework. The OBDHF formulation is expected to yield improved consistency between the energy and the one-particle density matrix relative to conventional double-hybrid functionals, with potential benefits for the computation of electron densities, dipole moments, and response properties. A comprehensive numerical implementation of the OBDHF equations, together with an extensive assessment of its performance across thermochemistry, kinetics, non-covalent interactions, and molecular properties benchmarks, is currently underway.